OpenStudy (anonymous):

Find the simplified form of the expression. Give your answer in scientific notation.

OpenStudy (anonymous):

\[(9 \times 10^{5})( 6 \times 10^{-7})\]

OpenStudy (anonymous):

Do you know how to do this @UnkleRhaukus

OpenStudy (unklerhaukus):

yes

OpenStudy (unklerhaukus):

\[(9 \times 10^{5})( 6 \times 10^{-7})\]\[=(9 \times6 )( 10^{5} \times 10^{-7})\]\[=(9 \times6 )( 10^{5-7})\]\[=\]

OpenStudy (unklerhaukus):

also, remember that scientific notation has exactly 1 numeral before the decimal place,

OpenStudy (anonymous):

then its 54x10^-2 = 54/100 =0.54

OpenStudy (anonymous):

5.4x10^-1

OpenStudy (anonymous):

would be the notation required i hope....

OpenStudy (anonymous):

Yes it is and that is one of my choices @sriramkumar Thanks, I think Im understanding it now

OpenStudy (unklerhaukus):

\[=(5.4)( 10^{5-7+1})\]